OpenStudy (anonymous):

If a coin is flipped 10 times, what is the probability that it will show all heads or all tails?

OpenStudy (anonymous):

Solving, one moment. ;)

OpenStudy (anonymous):

Are there choices?

OpenStudy (anonymous):

yes there is :) 1 - 2 1 --- - 512 1 -- - - 1,022 1 - - - - 1,024 1 ---- - 2,048

OpenStudy (anonymous):

i was thinking /1,022

OpenStudy (anonymous):

1/0,22*

OpenStudy (anonymous):

I'm not quite sure...ask AccessDenied. ;) Sorry...

OpenStudy (anonymous):

its okay , thanks for trying to help though :)

OpenStudy (anonymous):

No Problamo! ;)

OpenStudy (anonymous):

@AccessDenied He may need your help...not sure. :]

OpenStudy (accessdenied):

We have to find the probability of getting all heads, and of getting all tails individually. The probably of getting either is just the addition of the two probabilities, because both are favorable outcomes of all of the possible outcomes we could get. That part is clear?

OpenStudy (accessdenied):

So flipping the coin gives us a total number of possibilities of 2^10, 2 for every coin flip and multiplied together so that we are counting all the different choices of either heads or tails. There is only one way to flip all heads, and there is only one way to flip all tails... so we have 1/2^10 for each. 1/2^10 + 1/2^10 = 2/2^10 = 1/2^9